A truck traveling at a velocity of 33 m/s due east comes to a halt by decelerating at 11 m/s2. How far has the truck traveled while in the process of stopping?

 

 

 



1
Expert's answer
2021-09-23T08:32:58-0400
"v^2=v_0^2+2ad,""0=v_0^2+2ad,""d=-\dfrac{v_0^2}{2a}=-\dfrac{(33\ \dfrac{m}{s})^2}{2\cdot(-11\ \dfrac{m}{s^2})}=49.5\ m."